# GNN-DT: Graph Neural Network Enhanced Decision Transformer for Efficient Optimization in Dynamic Environments

### v1

@@ Abstract
Reinforcement Learning (RL) methods used for solving real-world optimization problems often involve dynamic state-action spaces, larger scale, and sparse rewards, leading to significant challenges in convergence, scalability, and efficient exploration of the solution space. This study introduces GNN-DT, a novel Decision Transformer (DT) architecture that integrates Graph Neural Network (GNN) embedders with a novel residual connection between input and output tokens crucial for handling dynamic environments. By learning from previously collected trajectories, GNN-DT reduces dependence on accurate simulators and tackles the sparse rewards limitations of online RL algorithms. We evaluate GNN-DT on the complex electric vehicle (EV) charging optimization problem and prove that its performance is superior and requires significantly fewer training trajectories, thus improving sample efficiency compared to existing DT baselines. Furthermore, GNN-DT exhibits robust generalization to unseen environments and larger action spaces, addressing a critical gap in prior DT-based approaches.
Electric Vehicle Charging, Graph Neural Networks, Large Language Models, Mathematical Optimization, Offline Reinforcement Learning
@@ 1 Introduction
Sequential decision-making problems are critical for efficiently operating a wide array of industries, such as power systems control (Roald et al., 2023), logistics optimization (Konstantakopoulos et al., 2022), portfolio management (Gunjan & Bhattacharyya, 2023), and advanced manufacturing processes (Gupta & Gupta, 2020). However, many practical problems, such as the electric vehicle (EV) charging optimization (Panda & Tindemans, 2024), are large-scale, have temporal dependencies, and aggregated constraints, often making conventional methods impractical (Bubeck, 2015). This is especially observed in dynamic environments, where the optimization landscape continuously evolves, requiring real-time solutions.
@@
Reinforcement learning (RL) (Sutton & Barto, 2018) has been extensively studied for solving optimization problems due to its ability to manage uncertainty, adapt to dynamic environments, and enhance decision-making through trial-and-error (Lan et al., 2023; Zhang et al., 2023a). In complex and large-scale scenarios, RL can provide high-quality solutions in real-time compared to traditional mathematical programming techniques that fail to do so (Jaimungal, 2022). However, RL approaches face significant challenges, such as sparse reward signals that slow learning and hinder convergence to optimal policies (Dulac-Arnold et al., 2021). In addition, RL solutions struggle to generalize when deployed in environments different from the one they were trained in, limiting their applicability in real-world scenarios with constantly changing conditions (Wang, 2024).
@@
Decision Transformers (DT) (Chen et al., 2021) is an offline RL algorithm that reframes traditional RL problems as generative sequence modeling tasks conditioned on future rewards (Zhang et al., 2023c). By learning from historical data, DTs effectively address the sparse reward issue inherent in online RL, relying on demonstrated successful outcomes instead of extensive trial-and-error exploration. However, the trajectory-stitching mechanism of DT often proves insufficient in dynamic real-world environments, leading to suboptimal policies. Although improved variants such as Q-regularized DT (Q-DT) (Hu et al., 2024) incorporate additional constraints for greater robustness, they still face significant challenges in generalizing across non-stationary tasks (Paster et al., 2024). Consequently, further architectural advances and training strategies are essential to ensure consistent performance in complex environments.
@@
This study introduces GNN-DT [^1] The code can be found at https://github.com/StavrosOrf/DT4EVs., a novel DT architecture that leverages the permutation-equivariant properties of Graph Neural Networks (GNNs) to handle dynamically changing state-action spaces (i.e., varying numbers of nodes over time) and improve generalization. By generating embeddings that remain consistent under node reordering, GNNs offer a powerful way to capture relational information in complex dynamic environments. Moreover, GNN-DT features a novel residual connection between input and output tokens, ensuring that action outputs are informed by the dynamically learned state embeddings for more robust decision-making. To demonstrate the superior performance of the proposed method, we conduct extensive experiments on the complex multi-objective EV charging optimization problem (Orfanoudakis et al., 2024a), which encompasses sparse rewards, temporal dependencies, and aggregated constraints. The main contributions are summarized as follows:
@@
  • Introducing a novel DT architecture integrating GNN embeddings, resulting in enhanced sample efficiency, superior performance, robust generalization to unseen environments, and effective scalability to larger action spaces, demonstrating the critical role of GNN-based embeddings in the model’s improvement.
@@
  • Demonstrating that online RL algorithms and offline DT baselines, even when trained on diverse datasets (Optimal, Random, Business-as-Usual) with varying sample sizes, perform inferior to GNN-DT when dealing with real-world optimization tasks.
  • Proving that both the size and type of training dataset critically influence the learning process of DTs, highlighting the importance of dataset selection.
@@
  • Highlighting that strategically integrating high- and low-quality training data (Optimal & Random datasets) significantly enhances policy learning, outperforming models trained exclusively on single-policy datasets.
@@ 2 Related Work
Classic DT encounters significant challenges, including limited trajectory stitching capabilities and difficulties in adapting to online environments. To address these issues, several enhancements have been proposed. The Q-DT (Hu et al., 2024) improves the ability to derive optimal policies from sub-optimal trajectories by relabeling return-to-go values in the training data. Elastic DT (Wu et al., 2024) enhances classic DT by enabling trajectory stitching during action inference at test time, while Multi-Game DT (Lee et al., 2024) advances its task generalization capabilities. The Online DT (Zheng et al., 2022; Villarrubia-Martin et al., 2023) extends DTs to online settings by combining offline pretraining with online fine-tuning, facilitating continuous policy updates in dynamic environments. Additionally, adaptations for offline safe RL incorporate cost tokens alongside rewards (Liu et al., 2023; Hong et al., 2024). DT has also been effectively applied to real-world domains, such as healthcare (Zhang et al., 2023c) and chip design (Lai et al., 2023), showcasing its versatility and practical utility.
@@
RL algorithms offer notable advantages for EV dispatch, including the ability to handle nonlinear models, robustly quantify uncertainty, and deliver faster computation than traditional mathematical programming (Qiu et al., 2023). Popular methods, such as DDPG (Jin et al., 2022), SAC (Jin & Xu, 2021), and batch RL (Sadeghianpourhamami et al., 2020), show promise but often lack formal constraint satisfaction guarantees and struggle to scale with high-dimensional state-action spaces (Yılmaz et al., 2024; Li et al., 2022). Safe RL frameworks address these drawbacks by imposing constraints via constrained MDPs, but typically sacrifice performance and scalability (Zhang et al., 2023b; Chen & Shi, 2022). Multiagent RL techniques distribute complexity across multiple agents, e.g. charging points, stations, or aggregators (Kamrani et al., 2025), yet still encounter convergence challenges and may underperform in large-scale applications. To the best of our knowledge, no study has used DTs for solving the complex EV charging problem, despite DT’s potential to handle sparse rewards effectively.
@@ 3.1 Offline RL
Offline RL aims to learn a policy $\pi_{\theta}(a\!\mid\!s)$ that maximizes the expected discounted return $\mathbb{E}\bigl{[}\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\bigr{]}$ without additional interactions with the environment (Levine et al., 2020). A Markov Decision Process (MDP) is defined by the tuple $(S,A,P,R,\gamma)$, where $S$ is the state space, $A$ the action space, $P$ the transition function, $R$ the reward function, $\gamma\in(0,1]$ the discount factor (Sutton & Barto, 2018). In the offline setting, a static dataset $\mathcal{D}=\{(s,a,r)\}$, collected by a (potentially suboptimal) policy, is provided. DTs leverage this dataset by treating RL trajectories as sequences, learning to predict actions that maximize returns based on previously collected experiences. A key component in DTs is the return-to-go (RTG), which for a time step $t$ can be defined as:
$G_{t}=\sum_{\tau=t}^{T}\gamma^{\tau-t}\,r_{\tau},$ (1)
representing the discounted cumulative reward from $t$ until the terminal time $T$. This formulation is particularly beneficial when real-time exploration is costly or impractical, while sufficient historical data remain available for training.
@@ 3.2 The EV Smart Charging Problem
We consider a set of $\mathcal{I}$ charging stations indexed i, all assumed to be controlled by a charge point operator (CPO) over a time window $\mathcal{T}$, divided into $T$ non-overlapping length intervals $\Delta t>0$. For a given time window, each charging station $i$ operates a set of $\mathcal{J}$ non-overlapping charging sessions, denoted by $\mathcal{J}_{i}=\{j_{i,1},\cdots,j_{i,J_{i}}\}$, where $J_{i,j}$ represents the $j^{th}$ charging event at the $i^{th}$ charging station and $J_{i}=|\mathcal{J}_{i}|$ is the total number of charging sessions seen by charging station i in an episode. A charging session is then represented as $j_{j,c}:\{t^{a}_{j,i},t^{d}_{j,i},\bar{p}_{j,i},e^{*}_{j,i}\},~{}\forall j,i$, where $t^{a},t^{d},\bar{p}$ and $e^{*}$ represent the arrival time, departure time, maximum charging power, and the desired battery energy level at the departure time. The primary goal is to minimize the total energy cost given by:
$f_{1}(p^{+},p^{-})=\sum_{t\in\mathcal{T}}\sum_{i\in\mathcal{I}}\Delta t(\Pi^{+}_{t}p_{i,t}^{+}-\Pi^{-}_{t}p_{i,t}^{-})$ (2)
$p^{+}_{i,t}$ and $p^{-}_{i,t}$ denote the charging or discharging power of the $i^{th}$ charging station during time interval $t$. $\Pi^{+}_{t}$ and $\Pi^{-}_{t}$ are the charging and discharging costs, respectively. Along with minimizing the total energy costs , the CPO also wants the aggregate power of all the charging stations ($p^{\sum}_{t}=\sum_{i\in\mathcal{I}}p^{+}_{i,t}-p^{-}_{i,t}$) to remain below the set power limit $p_{t}^{*}$. By doing so, the CPO avoids paying penalties due to overuse of network capacity. As the set point keeps on updated based on external factors, we introduce the penalty:
$f_{2}(p^{+},p^{-})=\sum_{t\in\mathcal{T}}\max\{0,p_{t}^{\sum}-p^{*}_{t}\},$ (3)
Maintaining the desired battery charge at departure is important for EV user satisfaction, which we model as:
$f_{3}(p^{+},p^{-})=\sum_{i\in\mathcal{I}}\sum_{j\in\mathcal{J}_{i}}\left(\sum_{t=t_{j,i}^{a}}^{{t_{j,i}^{a}}}(p_{i,t}^{+}-p_{i,t}^{-})-e^{*}_{j,i}\right)^{2}$ (4)
Eq. (4) defines a sparse reward added at each EV departure based on its departure energy level. Building on these objectives, the EV charging problem is formulated as:
@@
The multi-objective optimization function in Eq.5 integrates Eqs.24 using experimentally determined coefficients based on practical importance. This mixed integer programming problem is subject to lower-level operational constraints (e.g., EV battery, power levels) as detailed in AppendixA.1.
@@
The optimal EV charging problem can be framed as an MDP: $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},R)$. At any time step $t$, the state $\boldsymbol{s}_{t}\in\mathcal{S}$ is represented by a graph $\mathcal{G}_{t}=(\mathcal{N}_{t},\mathcal{E}_{t})$, where $\mathcal{N}_{t}$ is the set of nodes and $\mathcal{E}_{t}$ is the set of edges. Each node $n\in\mathcal{N}_{t}$ has a feature vector $\boldsymbol{x}_{n}^{t}\in\mathbb{R}^{d}$, capturing node-dependent information such as power limits and prices. This graph structure (Orfanoudakis et al., 2024b) efficiently models evolving relationships among EVs, chargers, and the grid infrastructure. The action space $\mathbf{a}_{t}\in\mathcal{A}$ is represented by a dynamic [^2] Note that the number of nodes in the state and action graph can vary in each step, because EVs arrive and depart. graph $\mathcal{G}^{\mathbf{a}}_{t}=(\mathcal{N}^{\mathbf{a}}_{t},\mathcal{E}^{\mathbf{a}}_{t})$, where nodes $\mathcal{N}^{\mathbf{a}}_{t}$ correspond to the decision variables of the optimization problem (e.g., EVs). Each node $n\in\mathcal{N}^{\mathbf{a}}_{t}$ represents a single action $a_{i,t}\in\mathbf{a}_{t}$, scaled by the corresponding charger’s maximum power limit. For charging, $a_{i,t}\in[0,1]$, and for discharging, $a_{i,t}\in[-1,0)$. The transition function $\mathcal{P}(\boldsymbol{s}_{t+1}\mid\boldsymbol{s}_{t},\mathbf{a}_{t})$ accounts for uncertainties in EV arrivals, departures, energy demands, and grid fluctuations. Finally, the reward $R(\boldsymbol{s}_{t},\mathbf{a}_{t})$ aligns with the objective described in Eq. 5, guiding the policy to maximize cost savings, respect operational constraints, and meet EV driver requirements. While Eqs.2 and 3 represent individual EV rewards and aggregated EV penalties, respectively, Eq.4 introduces a sparse reward that activates only when an EV departs, thereby creating complex temporal dependencies.

### v2

# GNN-DT: A Graph Neural Network Enhanced Decision Transformer for Efficient Optimization in Dynamic Environments

@@ Abstract
Reinforcement Learning (RL) methods used for solving real-world optimization problems often involve dynamic state-action spaces, larger scale, and sparse rewards, leading to significant challenges in convergence, scalability, and efficient exploration of the solution space. This study introduces GNN-DT, a novel Decision Transformer (DT) architecture that integrates Graph Neural Network (GNN) embedders with a novel residual connection between input and output tokens crucial for handling dynamic environments. By learning from previously collected trajectories, GNN-DT tackles the sparse rewards limitations of online RL algorithms and delivers high-quality solutions in real-time. We evaluate GNN-DT on the complex electric vehicle (EV) charging optimization problem and prove that its performance is superior and requires significantly fewer training trajectories, thus improving sample efficiency compared to existing DT and offline RL baselines. Furthermore, GNN-DT exhibits robust generalization to unseen environments and larger action spaces, addressing a critical gap in prior offline and online RL approaches.
@@ 1 Introduction
Sequential decision-making problems are critical for efficiently operating a wide array of industries, such as power systems control [1], logistics optimization [2], portfolio management [3], and advanced manufacturing processes [4]. However, many practical problems, such as the electric vehicle (EV) charging optimization [5], are large-scale, have temporal dependencies, and aggregated constraints, often making conventional methods impractical [6]. This is especially observed in dynamic environments, where the optimization landscape continuously evolves, requiring real-time solutions.
@@
Reinforcement learning (RL) [7] has been extensively studied for solving optimization problems due to its ability to manage uncertainty, adapt to dynamic environments, and enhance decision-making through trial-and-error [8, 9]. In complex and large-scale scenarios, RL can provide high-quality solutions in real-time compared to traditional mathematical programming techniques that fail to do so [10]. However, RL approaches face significant challenges, such as sparse reward signals that slow learning and hinder convergence to optimal policies [11]. In addition, RL solutions struggle to generalize when deployed in environments different from the one they were trained in, limiting their applicability in real-world scenarios with constantly changing conditions [12].
@@
Decision Transformers (DT) [13] is an offline RL algorithm that reframes traditional RL problems as generative sequence modeling tasks conditioned on future rewards [14]. By learning from historical data, DTs effectively address the sparse reward issue inherent in online RL, relying on demonstrated successful outcomes instead of extensive trial-and-error exploration. However, the trajectory-stitching mechanism of DT often proves insufficient in dynamic real-world environments, leading to suboptimal policies. Although improved variants such as Q-regularized DT (Q-DT) [15] incorporate additional constraints for greater robustness, they still face significant challenges in generalizing across non-stationary tasks [16]. Consequently, further architectural advances and training strategies are essential to ensure consistent performance in complex environments.
@@
This study introduces GNN-DT [^1] The code can be found at https://github.com/StavrosOrf/DT4EVs. , a novel DT architecture that leverages the permutation-equivariant properties of Graph Neural Networks (GNNs) to handle dynamically changing state-action spaces (i.e., varying numbers of nodes over time) and improve generalization. By generating embeddings that remain consistent under node reordering, GNNs offer a powerful way to capture relational information in complex dynamic environments. Moreover, GNN-DT features a novel residual connection between input and output tokens, ensuring that action outputs are informed by the dynamically learned state embeddings for more robust decision-making. To demonstrate the superior performance of the proposed method, we conduct extensive experiments on the complex multi-objective EV charging optimization problem [17], which encompasses sparse rewards, temporal dependencies, and aggregated constraints. The main contributions are summarized as follows:
@@
  • Introducing a novel DT architecture integrating GNN embeddings, resulting in enhanced sample efficiency, superior performance, robust generalization to unseen environments, and effective scalability to larger action spaces, demonstrating the critical role of GNNs.
@@
  • Demonstrating that online and offline RL baselines, even when trained on diverse datasets (Optimal, Random, Business-as-Usual) with varying sample sizes, perform inferior to GNN-DT when dealing with real-world optimization tasks.
@@
  • Proving that both the size and type of training dataset critically influence the learning process of DTs, highlighting the importance of dataset selection. Also, shown that strategically integrating high- and low-quality training data (Optimal & Random datasets) significantly enhances policy learning, outperforming models trained exclusively on single-policy datasets.
@@ 2 Related Work
Classic DT encounters significant challenges, including limited trajectory stitching capabilities and difficulties in adapting to online environments. To address these issues, several enhancements have been proposed. The Q-DT [15] improves the ability to derive optimal policies from sub-optimal trajectories by relabeling return-to-go values in the training data. Elastic DT [18] enhances classic DT by enabling trajectory stitching during action inference at test time, while Multi-Game DT [19] advances its task generalization capabilities. The Online DT [20, 21] extends DTs to online settings by combining offline pretraining with online fine-tuning, facilitating continuous policy updates in dynamic environments. Additionally, adaptations for offline safe RL incorporate cost tokens alongside rewards [22, 23]. DT has also been effectively applied to real-world domains, such as healthcare [14] and chip design [24], showcasing its practical utility.
@@
RL algorithms offer notable advantages for EV dispatch, including the ability to handle nonlinear models, robustly quantify uncertainty, and deliver faster computation than traditional mathematical programming [25]. Popular methods, such as DDPG [26], SAC [27], and batch RL [28], show promise but often lack formal constraint satisfaction guarantees and struggle to scale with high-dimensional state-action spaces [29, 30]. Safe RL frameworks address these drawbacks by imposing constraints via constrained MDPs, but typically sacrifice performance and scalability [31, 32]. Multiagent RL techniques distribute complexity across multiple agents, e.g., charging points, stations, or aggregators [33], yet still encounter convergence challenges and may underperform in large-scale applications. To the best of our knowledge, no study has used DTs to solve the complex EV charging problem, despite DTs’ potential to handle sparse rewards effectively.
@@ 3.1 Offline RL
Offline RL aims to learn a policy $\pi_{\theta}(a\!\mid\!s)$ that maximizes the expected discounted return $\mathbb{E}\bigl{[}\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\bigr{]}$ without additional interactions with the environment [34]. A Markov Decision Process (MDP) is defined by the tuple $(S,A,P,R,\gamma)$, where $S$ is the state space, $A$ the action space, $P$ the transition function, $R$ the reward function, $\gamma\in(0,1]$ the discount factor [7]. In the offline setting, a static dataset $\mathcal{D}=\{(s,a,r)\}$, collected by a (potentially suboptimal) policy, is provided. DTs leverage this dataset by treating RL trajectories as sequences, learning to predict actions that maximize returns based on previously collected experiences. A key component in DTs is the return-to-go (RTG), which for a time step $t$ can be defined as: $G_{t}=\sum_{\tau=t}^{T}\gamma^{\tau-t}\,r_{\tau},$ representing the discounted cumulative reward from $t$ until the terminal time $T$. Offline RL is particularly beneficial when real-time exploration is costly or impractical, while sufficient historical data are available.
@@ 3.2 The EV Smart Charging Problem
Working closely with a charge point operator (CPO), it was evident that existing heuristic and mathematical programming charging strategies don’t scale efficiently as EV fleets grow. To address this, we designed the state–action space and objectives around real-world operational constraints and assumptions provided by the CPO. We consider a set of $\mathcal{I}$ charging stations indexed i, all assumed to be controlled by a CPO over a time window $\mathcal{T}$, divided into $T$ non-overlapping intervals. For a given time window, each charging station $i$ operates a set of $\mathcal{J}$ non-overlapping charging sessions, denoted by $\mathcal{J}_{i}=\{j_{1,i},\cdots,j_{J_{i},i}\}$, where $j_{j,i}$ represents the $j^{th}$ charging event at the $i^{th}$ charging station and $J_{i}=|\mathcal{J}_{i}|$ is the total number of charging sessions seen by charging station i in an episode. A charging session is then represented as $j_{j,i}:\{t^{a}_{j,i},t^{d}_{j,i},\bar{p}_{j,i},e^{*}_{j,i}\},~{}\forall j,i$, where $t^{a},t^{d},\bar{p}$ and $e^{*}$ represent the arrival time, departure time, maximum charging power, and the desired battery energy level at the departure time. The primary goal is to minimize the total energy cost given by:
$f_{1}(p^{+},p^{-})=\sum\nolimits_{t\in\mathcal{T}}\sum\nolimits_{i\in\mathcal{I}}\Delta t\bigl{(}\Pi^{+}_{t}\,p_{i,t}^{+}-\Pi^{-}_{t}\,p_{i,t}^{-}\bigr{)}$ (1)
$p^{+}_{i,t}$ and $p^{-}_{i,t}$ denote the charging or discharging power of the $i^{th}$ charging station during time interval $t$. $\Pi^{+}_{t}$ and $\Pi^{-}_{t}$ are the charging and discharging costs, respectively. Along with minimizing the total energy costs , the CPO also wants the aggregate power of all the charging stations ($p^{\sum}_{t}=\sum_{i\in\mathcal{I}}p^{+}_{i,t}-p^{-}_{i,t}$) to remain below the set power limit $p_{t}^{*}$. By doing so, the CPO avoids paying penalties due to overuse of network capacity. Hence, we introduce the penalty:
$f_{2}(p^{+},p^{-})=\sum\nolimits_{t\in\mathcal{T}}\max\{0,\,p_{t}^{\sum}-p^{*}_{t}\},$ (2)
Maintaining the desired battery charge at departure is crucial for EV user satisfaction. We model as:
$f_{3}(p^{+},p^{-})=\sum\nolimits_{i\in\mathcal{I}}\,\sum\nolimits_{j\in\mathcal{J}_{i}}\Bigl{(}\sum\nolimits_{t=t^{a}_{j,i}}^{\,t^{d}_{j,i}}(p_{i,t}^{+}-p_{i,t}^{-})-e^{*}_{j,i}\Bigr{)}^{2}$ (3)
Eq. (3) defines a sparse reward added at each EV departure based on its departure energy level. Building on these objectives, the overall EV charging problem is formulated as:
@@
The multi-objective optimization function in Eq.4 integrates Eqs.13 using experimentally determined coefficients based on practical importance. This mixed integer programming (MIP) problem is subject to lower-level operational constraints (e.g., EV battery, power levels) as detailed in AppendixA.1.
@@
The optimal EV charging problem can be framed as an MDP: $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},R)$. At any time step $t$, the state $\boldsymbol{s}_{t}\in\mathcal{S}$ is represented by a dynamic graph $\mathcal{G}_{t}=(\mathcal{N}_{t},\mathcal{E}_{t})$, where $\mathcal{N}_{t}$ is the set of nodes and $\mathcal{E}_{t}$ is the set of edges. The graph is dynamic since the number of nodes in the state and action graph can vary in each step, because of EVs’ arrival and departures. Each node $n\in\mathcal{N}_{t}$ has a feature vector $\boldsymbol{x}_{n}^{t}\in\mathbb{R}^{d}$, capturing node-dependent information such as power limits and prices (see Appendix A.3 for the complete definition). This graph structure [35] efficiently models evolving relationships among EVs, chargers, and the grid infrastructure. The action space $\mathbf{a}_{t}\in\mathcal{A}$ is represented by a dynamic graph $\mathcal{G}^{\mathbf{a}}_{t}=(\mathcal{N}^{\mathbf{a}}_{t},\mathcal{E}^{\mathbf{a}}_{t})$, where nodes $\mathcal{N}^{\mathbf{a}}_{t}$ correspond to the decision variables of the optimization problem (e.g., EVs). Each node $n\in\mathcal{N}^{\mathbf{a}}_{t}$ represents a single action $a_{i,t}\in\mathbf{a}_{t}$, scaled by the corresponding charger’s maximum power limit. For charging, $a_{i,t}\in[0,1]$, and for discharging, $a_{i,t}\in[-1,0)$. The transition function $\mathcal{P}(\boldsymbol{s}_{t+1}\mid\boldsymbol{s}_{t},\mathbf{a}_{t})$ accounts for uncertainties in EV arrivals, departures, energy demands, and grid fluctuations. Finally, the reward function is the same with the objective described in Eq. 4 and is defined as $R(\boldsymbol{s}_{t},\mathbf{a}_{t})=f_{1}-100f_{2}-10f_{3}$ for a single timestep ($\mathcal{T}=\{t\}$). The reward is guiding the policy to maximize cost savings, respect operational constraints, and meet EV driver requirements. While Eqs.1 and 2 represent individual EV rewards and aggregated EV penalties, respectively, Eq.3 introduces a sparse reward that activates only when an EV departs, thereby creating complex temporal dependencies. The reward function weights were selected after comprehensive experimentation, ensuring the optimal solution of the MIP remains the same while the RL training is balanced between exploration and penalization.